# Efficient Transformed Gaussian Process State-Space Models for Non-Stationary High-Dimensional Dynamical Systems

### v3

@@ Abstract
Gaussian process state-space models (GPSSMs) offer a principled framework for learning and inference in nonlinear dynamical systems with uncertainty quantification. However, existing GPSSMs are limited by the use of multiple independent stationary Gaussian processes (GPs), leading to prohibitive computational and parametric complexity in high-dimensional settings and restricted modeling capacity for non-stationary dynamics. To address these challenges, we propose an efficient transformed Gaussian process state-space model (ETGPSSM) for scalable and flexible modeling of high-dimensional, non-stationary dynamical systems. Specifically, our ETGPSSM integrates a single shared GP with input-dependent normalizing flows, yielding an expressive implicit process prior that captures complex, non-stationary transition dynamics while significantly reducing model complexity. For the inference of the implicit process, we develop a variational inference algorithm that jointly approximates the posterior over the underlying GP and the neural network parameters defining the normalizing flows. To avoid explicit variational parameterization of the latent states, we further incorporate the ensemble Kalman filter (EnKF) into the variational framework, enabling accurate and efficient state estimation. Extensive empirical evaluations on synthetic and real-world datasets demonstrate the superior performance of our ETGPSSM in system dynamics learning, high-dimensional state estimation, and time-series forecasting, outperforming existing GPSSMs and neural network-based SSMs in terms of computational efficiency and accuracy.
Index Terms:
Gaussian process, state-space model, normalizing flows, variational inference, high-dimensional dynamical systems.
@@ I Introduction
State-space models (SSMs) provide a flexible framework for modeling dynamical systems, where latent states evolve based on internal dynamics and external inputs [1]. When system dynamics are fully known, Bayesian filtering methods—such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF)—enable sequential state estimation by combining prior knowledge with incoming observations, yielding reliable predictions under uncertainty. This capability makes SSMs widely applicable in robotics, finance, and signal processing, where modeling temporal dependencies is crucial [2, 3, 4].
@@
However, in complex real-world systems, including nonlinear and high-dimensional problems such as climate modeling, physiological processes, and neural dynamics, the underlying dynamics may be too intricate or poorly understood to specify a physics-informed or explicitly parameterized model. In these cases, traditional filtering methods cannot be applied in their standard form and must instead rely on data-driven approaches to learn the dynamics from observations, complicating the task of accurate and robust state estimation [5].
@@
To address this challenge, a variety of data-driven system identification approaches have been proposed. Deterministic models, such as deep state-space models (DSSMs), employ neural networks to learn system dynamics and have gained popularity for their ability to capture high-dimensional, nonlinear behaviors [6]. However, these models often require large training datasets to avoid overfitting, and their generalization performance may degrade in uncertain or data-scarce environments. Moreover, their predictions can be difficult to interpret, especially when compared to more structured probabilistic models [7]. As a result, DSSMs may face challenges in practical deployment despite their modeling capacity [8].
@@
Stochastic models based on flexible random process priors provide a principled alternative within a Bayesian learning framework [9]. A prominent example is the class of Gaussian process state-space models (GPSSMs), which leverage Gaussian processes (GPs) to capture uncertainty and provide implicit regularization through prior structure. These properties make GPSSMs attractive for applications where uncertainty quantification and data efficiency are critical [10, 11]. GPSSMs have been applied in diverse domains, including human pose and motion learning [12], robotics and control [13], reinforcement learning [14, 15], target tracking and navigation [16], and magnetic-field sensing [17].
@@
Research on GPSSMs has also focused on advancing learning and inference methodologies. Early works either assumed a pre-learned GPSSM and concentrated solely on latent state inference—an impractical assumption for many real-world applications [18, 19, 20]—or employed maximum a posteriori (MAP) estimation to jointly infer the model and the latent states [12, 21], which is prone to overfitting. To address these issues, a important development came with the fully Bayesian treatment of GPSSMs using particle Markov chain Monte Carlo (PMCMC) methods [22], though the inherent cubic computational complexity of GP remains prohibitive for long time series. Subsequent efforts [23, 24, 25, 26] adopted reduced-rank GP approximations to alleviate the computational demands, yet PMCMC scalability remains a challenge. This has led to a growing interest in variational inference methods [10, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] that can leverage sparse GP approximations with inducing points to enhance scalability [40, 41].
@@
However, existing GPSSMs, which employ multiple independent GPs to model system dynamics, remain limited to low-dimensional state spaces and face fundamental challenges when scaling to high-dimensional systems. Specifically, in variational GPSSMs, each output dimension of the state transition function is modeled by a separate GP, leading to a total computational cost that scales linearly with the state dimensionality. Since each GP has a cubic complexity in the number of data points due to covariance matrix inversion, the overall computational burden becomes prohibitive. Moreover, the number of model parameters grows quadratically with the state dimension, further exacerbating scalability issues. Additionally, all GPSSMs rely on stationary GP priors, which assume that the statistical properties of the transition function are invariant under input translation. This assumption often fails for systems with time-varying or non-stationary dynamics, limiting the ability of GPSSMs to accurately capture complex real-world behaviors.
@@
To address these challenges, we propose a novel approach for learning and inference in GPSSMs tailored to non-stationary, high-dimensional dynamical systems. Our primary contribution is the development and integration of an efficient yet flexible implicit process prior [42, 43] for the state transition function. This prior combines the strengths of GPs, normalizing flows [44], and (Bayesian) neural networks, significantly enhancing modeling capabilities in high-dimensional state spaces. Our key contributions are summarized as follows:
@@
  • Instead of employing multiple independent GPs for the transition function, we propose applying an efficient transformed Gaussian process (ETGP) [42], an implicit random process that combines a shared GP with normalizing flows. Crucially, the flow parameters are made input-dependent through neural network modeling. This novel modeling approach offers two key advantages: (1) it captures complex, non-stationary transition functions that standard GP priors cannot represent, and (2) by utilizing only a single shared GP, it mitigates the computational and parametric complexities of traditional GPSSMs in high-dimensional spaces, enabling effective scaling to higher dimensions.
@@
Part of this work was presented at IEEE ICASSP 2024 [38]. The current paper introduces input-dependent normalizing flows to further enhance model flexibility and proposes an EnKF-aided variational inference algorithm for efficient state estimation. More extensive empirical validations are also included to demonstrate the superiority of the proposed method.
@@ II-A State-Space Models (SSMs)
The latent states evolve according to a first-order Markov process, meaning that at any time step $t\in\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ depends only on the current state ${\mathbf{x}}_{t}$ through the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$. In this work, the emission function is assumed to be linear, parameterized by a matrix $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$. For systems with nonlinear emissions, the latent state can be augmented to a higher-dimensional representation, effectively rendering the emission function linear [30]. This reparameterization helps mitigate non-identifiability issues frequently encountered in data-driven SSMs [10]. Both latent states and observations are perturbed by zero-mean Gaussian noise, with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
@@ II-B Gaussian Process State-Space Models (GPSSMs)
As discussed in Section I, system dynamics in real-world complex scenarios are often unknown, with only a set of noisy observations, ${\mathbf{y}}_{1:T}\!=\!\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, available. Consequently, there has been growing interest in adopting data-driven approaches, such as GPs [9], to model these dynamics and quantify the underlying uncertainties.
@@ II-B1 Gaussian process (GP)
The GP is a generalization of the Gaussian distribution over infinite index sets, enabling the specification of distributions over functions $\tilde{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$. A GP is fully characterized by its mean function, often set to zero, and its covariance (kernel) function $k({\mathbf{x}},{\mathbf{x}}^{\prime})$, which includes a set of hyperparameters $\bm{\theta}_{gp}$ optimized during model selection [9]. By definition, the function values $\tilde{{\mathbf{f}}}=\{\tilde{f}({\mathbf{x}}_{i})\}_{i=1}^{N}\in\mathbb{R}^{N}$, evaluated at a finite set of inputs $\mathbf{X}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{x}}$, follow a joint Gaussian distribution:
@@
where $\mathbf{K}\in\mathbb{R}^{N\times N}$ is the covariance matrix with entries $[\mathbf{K}]_{i,j}=k({\mathbf{x}}_{i},{\mathbf{x}}_{j})$.
Given observed function values $\tilde{{\mathbf{f}}}$ at inputs $\mathbf{X}$, the GP prediction at a new input ${\mathbf{x}}_{*}$, denoted $p(\tilde{f}({\mathbf{x}}_{*})\mid\tilde{{\mathbf{f}}},\mathbf{X})$, is also Gaussian, with posterior mean $\xi({\mathbf{x}}_{*})$ and posterior variance $\Xi({\mathbf{x}}_{*})$ given by:
@@ II-B2 Gaussian process state-space model
Placing a GP prior over the transition function $\bm{f}(\cdot)$ in the classic SSM (see Eq. (1)) yields the GPSSM [10]:
@@
where the initial state $p({\mathbf{x}}_{0})$ is assumed known and Gaussian. A graphical representation is shown in Fig. 1.
When the state dimension $d_{x}>1$, the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ is typically modeled using $d_{x}$ independent GPs. Each component function $f_{d}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}$ maps to one output dimension:
@@
where each GP has its own kernel function and hyperparameters [30]. The central challenge in GPSSMs lies in jointly learning the transition dynamics and noise parameters, i.e., learning $[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}]$, while simultaneously inferring both the latent states and the underlying functions.
${{\mathbf{x}}}_{0}$ ${{\mathbf{x}}}_{1}$cdots $\!\!{{\mathbf{x}}}_{t-1}\!\!$ ${{\mathbf{x}}}_{t}$cdots $\cdots$ ${{\mathbf{f}}}_{1}$cdots $\!\!{{\mathbf{f}}}_{t-1}\!\!$ ${{\mathbf{f}}}_{t}$cdots ${{\mathbf{y}}}_{1}$cdots $\!\!{{\mathbf{y}}}_{t-1}\!\!$ ${{\mathbf{y}}}_{t}$cdots
@@ II-C Challenges in Existing Variational GPSSMs
Note that in Eq. (3), evaluating the GP posterior requires inverting the covariance matrix $\mathbf{K}$, incurring a computational cost of $\mathcal{O}(N^{3})$, which becomes prohibitive for large $N$. To address this, variational GPSSMs typically employ sparse GP approximations [40, 41], a widely used technique in GP-based models [45, 10] due to its analytical tractability and reduced computational cost. The core idea is to augment the GP with a smaller set of inducing points, which act as a sparse approximation of the full GP.
More specifically, we can introduce $M\!\ll\!N$ inducing inputs $\mathbf{Z}=\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ (where ${\mathbf{z}}_{i}\in\mathbb{R}^{d_{x}}$) and their corresponding inducing outputs ${\mathbf{u}}=\{\mathrm{u}_{i}\}_{i=1}^{M}=\{\tilde{f}({\mathbf{z}}_{i})\}_{i=1}^{M}$, which inherit the same GP prior as the original process, i.e., ${\mathbf{u}}\sim\mathcal{N}(\bm{0},\mathbf{K}_{\mathbf{Z},\mathbf{Z}})$. The augmented GP prior can be written as: [^1] With slight notation abuse, we denote the distribution of GP prior as $p(\tilde{f})$.
@@
To approximate the posterior, a variational distribution is defined over the $M$ inducing outputs, $q({\mathbf{u}})=\mathcal{N}({\mathbf{u}}\mid\mathbf{m},\mathbf{S})$, while the prior conditional is used for the rest of the GP:
@@
  • The variational approximation introduces a number of variational parameters that scale quadratically with the state dimensionality $d_{x}$. Specifically, variational GPSSMs define inducing inputs $\{\mathbf{Z}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times d_{x})}$, variational means $\{\mathbf{m}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times M}$, and covariance matrices $\{\mathbf{S}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times M)}$ for the $d_{x}$-dimensional inducing outputs $\{\mathbf{u}_{d}\}_{d=1}^{d_{x}}$. As the state dimension increases, the quadratic scaling of inducing inputs and the associated variational parameters leads to significant computational overhead [38].
  • Most GPSSMs impose a stationary GP prior over the transition function, as shown in Eq. (4a), to model time-invariant dynamics. However, such stationary assumptions may fail to capture time-varying or context-dependent behavior [^2] Here, “stationarity” refers to the GP prior over the transition function, not the observed time series. A GPSSM with a stationary kernel may still model non-stationary observed data.. Additionally, modeling each output dimension with an independent GP ignores correlations across dimensions, potentially limiting modeling capacity.
In the next section, we detail our proposed method to resolve these issues.

### v4

@@ Abstract
Gaussian process state-space models (GPSSMs) offer a principled framework for learning and inference in nonlinear dynamical systems with uncertainty quantification. However, existing GPSSMs are limited by the use of multiple independent stationary Gaussian processes (GPs), leading to prohibitive computational and parametric complexity in high-dimensional settings and restricted modeling capacity for non-stationary dynamics. To address these challenges, we propose an efficient transformed Gaussian process state-space model (ETGPSSM) for scalable and flexible modeling of high-dimensional, non-stationary dynamical systems. Specifically, our ETGPSSM integrates a single shared GP with input-dependent normalizing flows, yielding an expressive non-stationary implicit process prior that can capture complex transition dynamics while significantly reducing model complexity. For the inference of the implicit process, we develop a variational inference algorithm that jointly approximates the posterior over the underlying GP and the neural network parameters defining the normalizing flows. To avoid explicit variational parameterization of the latent states, we further incorporate the ensemble Kalman filter (EnKF) into the variational framework, enabling accurate and efficient state estimation. Extensive empirical evaluations on synthetic and real-world datasets demonstrate the superior performance of our ETGPSSM in system dynamics learning, high-dimensional state estimation, and time-series forecasting, outperforming existing GPSSMs and neural network-based SSMs in terms of computational efficiency and accuracy.
@@ I Introduction
State-space models (SSMs) provide a flexible framework for modeling dynamical systems, where latent states evolve based on internal dynamics and external inputs [sarkka2013bayesian]. When system dynamics are fully known, Bayesian filtering methods—such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF)—enable sequential state estimation by combining prior knowledge with incoming observations, yielding reliable predictions under uncertainty. This capability makes SSMs widely applicable in robotics, finance, and signal processing, where modeling temporal dependencies is crucial [kullberg2021online, khan2008distributing, tobar2015unsupervised].
@@
However, in complex real-world systems, such as climate modeling, physiological processes, and neural dynamics, the underlying dynamics are often highly nonlinear, high-dimensional, and insufficiently understood to permit a physics-informed or explicitly parameterized model. In such settings, traditional filtering methods cannot be applied in their standard form and must instead resort to data-driven approaches that learn the system dynamics directly from observations. This shift significantly complicates the task of achieving accurate and robust state estimation [revach2022kalmannet].
@@
A variety of data-driven approaches for modeling dynamical systems have been proposed. Deterministic models, such as deep state-space models (DSSMs), employ neural networks to learn system dynamics and have gained popularity for their ability to capture high-dimensional, nonlinear behaviors [girin2021dynamical]. However, these models often require large training datasets to avoid overfitting, and their generalization performance may degrade in uncertain or data-scarce environments. Moreover, their predictions can be difficult to interpret, especially when compared to more structured probabilistic models [krishnan2017structured]. As a result, DSSMs may face challenges in practical deployment despite their modeling capacity [ghosh2024danse].
@@
Stochastic models based on flexible random process priors provide a principled alternative within a Bayesian learning framework [williams2006gaussian]. A prominent example is the class of Gaussian process state-space models (GPSSMs), which leverage Gaussian processes (GPs) to encapsulate uncertainty and provide implicit regularization through prior structure. These properties make GPSSMs particularly attractive for applications where uncertainty quantification and data efficiency are critical [frigola2015bayesian, suwandi2023sparsityaware]. GPSSMs have been applied in diverse domains, including human pose and motion learning [wang2007gaussian], robotics and control [deisenroth2013gaussian], reinforcement learning [arulkumaran2017deep, yan2020gaussian], target tracking and navigation [xie2020learning], and magnetic-field sensing [berntorp2023constrained].
@@
In parallel to their practical applications, significant research efforts have been dedicated to advancing the learning and inference algorithms that underpin GPSSMs. Early works either assumed a pre-learned GPSSM and concentrated solely on latent state inference—an impractical assumption for many real-world applications [ko2009gp, deisenroth2009analytic, deisenroth2011robust]—or employed maximum a posteriori (MAP) estimation to jointly infer the model and the latent states [wang2007gaussian, ko2011learning], which is prone to overfitting. To address these issues, a important development came with the fully Bayesian treatment of GPSSMs using particle Markov chain Monte Carlo (PMCMC) methods [frigola2013bayesian], though the inherent cubic computational complexity of GP remains prohibitive for long time series. Subsequent efforts adopted reduced-rank GP approximations to alleviate the computational demands [svensson2016computationally, svensson2017flexible, berntorp2021online, liu2023sequential], yet PMCMC scalability remains a challenge. This has led to a growing interest in variational inference methods [frigola2015bayesian, frigola2014variational, mchutchon2015nonlinear, eleftheriadis2017identification, doerr2018probabilistic, ialongo2019overcoming, curi2020structured, liu2020gpssm, lindinger2022laplace, fan2023free, lin2023towards, lin2023ensemble, lin2023towards_efficient, lin2022output] that can leverage sparse GP approximations with inducing points to enhance scalability [titsias2009variational, hensman2013gaussian].
@@
However, existing GPSSMs, which employ multiple independent GPs to model system dynamics, remain limited to low-dimensional state spaces and face fundamental challenges when scaling to high-dimensional systems. Specifically, in variational GPSSMs, each output dimension of the state transition function is modeled by a separate GP, leading to a total computational cost that scales linearly with the state dimensionality. Since each GP has a cubic complexity in the number of data points due to covariance matrix inversion, the overall computational burden becomes prohibitive. Moreover, the number of model parameters grows quadratically with the state dimension, further exacerbating scalability issues. Additionally, all GPSSMs rely on stationary GP priors, which assume that the statistical properties of the transition function are invariant under input translation. This assumption often fails for systems exhibiting time-varying or non-stationary dynamics, limiting the expressiveness and practical applicability of conventional GPSSMs.
@@
To this end, we propose a novel GPSSM framework for non-stationary, high-dimensional dynamical systems, whose primary contribution is an efficient, flexible implicit process prior for the state transition function. By unifying Gaussian processes, input-dependent normalizing flows [kobyzev2020normalizing], and (Bayesian) neural networks, this prior is inherently non-stationary while maintaining computational and parametric complexity that scales gracefully with state dimensionality. The key contributions are summarized as follows:
@@
  • Instead of employing multiple independent GPs for the transition function, we propose applying an efficient transformed Gaussian process (ETGP) [maronas2022efficient], an implicit random process that combines a shared GP with normalizing flows. Crucially, the flow parameters are made input-dependent through neural network modeling. This novel modeling approach offers two key advantages: (1) it captures complex, non-stationary transition functions that standard GP priors cannot represent, and (2) by utilizing only a single shared GP, it mitigates the computational and parametric complexities of traditional GPSSMs in high-dimensional spaces, enabling effective scaling to higher dimensions.
@@
Part of this work was presented at IEEE ICASSP 2024 [lin2023towards_efficient]. The current paper introduces input-dependent normalizing flows to further enhance model flexibility and proposes an EnKF-aided variational inference algorithm for efficient state estimation. More extensive empirical validations are also included to demonstrate the superiority of the proposed method.
@@ II-A State-Space Models (SSMs)
The latent states evolve according to a first-order Markov process, meaning that at any time step $t\in\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ depends only on the current state ${\mathbf{x}}_{t}$ through the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$. In this work, the emission function is assumed to be linear, parameterized by a matrix $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$. For systems with nonlinear emissions, the latent state can be augmented to a higher-dimensional representation, effectively rendering the emission function linear [doerr2018probabilistic]. This reparameterization helps mitigate non-identifiability issues frequently encountered in data-driven SSMs [frigola2015bayesian]. Both latent states and observations are perturbed by zero-mean Gaussian noise, with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
${{\mathbf{x}}}_{0}$ ${{\mathbf{x}}}_{1}$ $\cdots$ ${{\mathbf{x}}}_{t-1}\!\!$ ${{\mathbf{x}}}_{t}$ $\cdots$ $\cdots$ ${{\mathbf{f}}}_{1}$ $\cdots$ ${{\mathbf{f}}}_{t-1}\!\!$ ${{\mathbf{f}}}_{t}$ $\cdots$ ${{\mathbf{y}}}_{1}$ $\cdots$ ${{\mathbf{y}}}_{t-1}\!\!$ ${{\mathbf{y}}}_{t}$ $\cdots$
@@ II-B Gaussian Process State-Space Models (GPSSMs)
As discussed in Section I, system dynamics in real-world complex scenarios are often unknown, but only a set of noisy observations, ${\mathbf{y}}_{1:T}\!=\!\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, where $T\in\mathbb{Z}_{+}$, is available. Consequently, there has been growing interest in adopting data-driven approaches, such as GPs [williams2006gaussian], to model these dynamics and quantify the underlying uncertainties.
@@ II-B1 Gaussian process (GP)
The GP is a generalization of the Gaussian distribution over infinite index sets, enabling the specification of distributions over functions $\tilde{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$.
Definition 1.
A stochastic process $\{\tilde{f}(\mathbf{x}):\mathbf{x}\in\mathcal{X}\}$ indexed by a domain $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ is said to be a GP [williams2006gaussian] if, for every finite set of inputs $\mathbf{X}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d_{x}}$, the corresponding random vector $\tilde{{\mathbf{f}}}=\{\tilde{f}({\mathbf{x}}_{i})\}_{i=1}^{N}$ is jointly Gaussian. A GP is fully characterized by its mean function, often set to zero, and its covariance (kernel) function $k({\mathbf{x}},{\mathbf{x}}^{\prime})$, which includes a set of hyperparameters $\bm{\theta}_{gp}$ [williams2006gaussian]. We denote this as
$\tilde{f}({\mathbf{x}})\sim\mathcal{GP}\big(0,\,k({\mathbf{x}},{\mathbf{x}}^{\prime}\,;{{\bm{\theta}}}_{gp})\big).$
A widely used family of kernels is the class of stationary kernels [williams2006gaussian], which is formally defined as follows.
Definition 2.
A kernel $k$ is stationary if there exists a function $\kappa:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{x}}\to\mathbb{R}$, such that $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$ for all ${\mathbf{x}},{\mathbf{x}}^{\prime}\subset\cal X$.
This assumption encodes translation-invariant properties and is shared by common kernels such as the squared exponential (SE) and Matérn families [williams2006gaussian]. Together with a constant mean function, a stationary kernel defines a stationary GP, whose finite-dimensional distributions are invariant to translations of the input space. Specifically, for any inputs $\mathbf{X}$, the distribution
@@
where $\mathbf{K}\in\mathbb{R}^{N\times N}$ and $\mathbf{K}]_{i,j}=k({\mathbf{x}}_{i},{\mathbf{x}}_{j}\,;{{\bm{\theta}}}_{gp})$, remains unchanged under translation.
Given function values $\tilde{{\mathbf{f}}}$, the GP prediction at a new input ${\mathbf{x}}_{*}\!\subset\!\mathcal{X}$, denoted $p(\tilde{f}({\mathbf{x}}_{*})|\tilde{{\mathbf{f}}},\mathbf{X})$, is also Gaussian, with posterior mean $\xi({\mathbf{x}}_{*})$ and posterior variance $\Xi({\mathbf{x}}_{*})$ given by:
@@ II-B2 Gaussian process state-space model
Placing a GP prior over the transition function $\bm{f}(\cdot)$ in the classic SSM (see Eq. (1)) yields the GPSSM [frigola2015bayesian]:
@@
where the initial state $p({\mathbf{x}}_{0})$ is assumed known and Gaussian for simplicity. A graphical representation is shown in Fig. 1.
When the state dimension $d_{x}>1$, the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$ is typically modeled using $d_{x}$ independent GPs. Each component function $f_{d}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$ maps to one output dimension:
@@
where each GP has its own kernel function and hyperparameters [doerr2018probabilistic]. Given a single observed sequence $\{\mathbf{y}_{t}\}_{t=1}^{T}$, the central task in GPSSMs involves jointly learning the transition dynamics and noise parameters, i.e., learning ${{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}]$, while simultaneously inferring both the corresponding latent states and the underlying functions.
@@ II-C Challenges in Existing Variational GPSSMs
In contrast to the standard GP described in Eq. (3), where the GP posterior is computed directly from observed input-output pairs, inferring the GP in a GPSSM is more complex. The key challenge is that the inputs to the transition function are the latent states $\{\mathbf{x}_{t}\}_{t=0}^{T-1}$, which are themselves unobserved. Consequently, the GP posterior over the transition function must be inferred jointly with the latent state trajectory. This is typically done approximately via variational inference [frigola2014variational]. The variational GPSSMs employ sparse GP approximations [titsias2009variational, hensman2013gaussian], a widely used technique in GP-based models [maronas2021transforming, frigola2015bayesian] due to its analytical tractability and reduced computational cost. The core idea is to augment the GP with a smaller set of inducing points, which act as a sparse approximation of the full GP.
More specifically, we can introduce $M\!\ll\!N$ inducing inputs $\mathbf{Z}=\{{\mathbf{z}}_{i}\}_{i=1}^{M}\subset\mathbb{R}^{d_{x}}$ and their corresponding inducing outputs ${\mathbf{u}}=\{\mathrm{u}_{i}\}_{i=1}^{M}=\{\tilde{f}({\mathbf{z}}_{i})\}_{i=1}^{M}$, which inherit the same GP prior as the original process, i.e., ${\mathbf{u}}\sim\mathcal{N}(\bm{0},\mathbf{K}_{\mathbf{Z},\mathbf{Z}})$. The augmented GP prior can be written as: [^1] With slight notation abuse, we denote the distribution of GP prior as p(f~)p($\tilde${f}).
@@
To approximate the posterior, a variational distribution is defined over the $M$ inducing outputs, $q({\mathbf{u}})\!=\!\mathcal{N}({\mathbf{u}}|\mathbf{m},\mathbf{S})$, while the prior conditional is used for the rest of the GP:
@@
  • The variational approximation introduces a number of variational parameters that scale quadratically with the state dimensionality $d_{x}$. Specifically, variational GPSSMs define inducing inputs $\{\mathbf{Z}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times d_{x})}$, variational means $\{\mathbf{m}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times M}$, and covariance matrices $\{\mathbf{S}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times M)}$ for the $d_{x}$-dimensional inducing outputs $\{\mathbf{u}_{d}\}_{d=1}^{d_{x}}$. As the state dimension increases, the quadratic scaling of inducing inputs and the associated variational parameters leads to significant computational overhead [lin2023towards_efficient].
  • Most GPSSMs impose a stationary GP prior on the transition function, as shown in Eq. (4a). However, this stationarity assumption may be inadequate for capturing time-varying or input-dependent dynamical behavior. Additionally, modeling each output dimension with an independent GP ignores potential correlations across state dimensions, which can limit the model’s capacity.
It is important to clarify that the “stationarity” in this work refers specifically to the properties of the transition function prior, not the observed time series $\mathbf{y}_{1:T}$. A non-stationary transition function is one whose characteristics (e.g., amplitude, smoothness, or periodicity) vary with the input state. Notably, a GPSSM with a stationary kernel in its transition function can still generate a non-stationary time series. In the next section, we present a methodology to resolve these issues.